OpenStudy (anonymous):

A. Find the remainder when 6x^1000 - 17x^562 + 12x +26 is divided by x+1. B. Is x-1 a factor of x^567 - 3x^400 + x^9 +2? I just need to see that the answer I have is correct :) thanks

OpenStudy (anonymous):

Tell me what you got as remainder??

OpenStudy (anonymous):

a. I got 27 using synthetic division.

OpenStudy (anonymous):

Well I am not quite sure.. In B part, to check x - 1 is the factor : Put x = 1 in the given expression and tell me after solving does it become 0 or not?

OpenStudy (anonymous):

I have told you.. If (x-1) is the factor then it must satisfy the given expression.. Plug in x = 1 in the given expression and check whether you get 0 or something else..

OpenStudy (callisto):

For a, are you sure you would like to use synthetic division? You'll have to write 998 '0's for that...

OpenStudy (callisto):

*997

OpenStudy (anonymous):

okay, then tell me another way, i'm obviously learning here.

OpenStudy (callisto):

Use remainder theorem... Let f(x) = 6x^1000 - 17x^562 + 12x +26 Since it is divided by x+1 remainder = f(-1) = 6(-1)^1000 - 17(-1)^562 + 12(-1) +26 =...

OpenStudy (anonymous):

okay thankyou

OpenStudy (callisto):

Welcome. You can use this to check of (x-a) is a factor of polynomial. If f(a) = 0, that means the remainder is 0, indicating that (x-a) is a factor of the polynomial.